Wildcat Electronics sells audio and video equipment. Its billing division has information on the age of its accounts receivables. The data collected indicate that the age of the accounts follows a normal distribution with a mean of 12 days and a standard deviation of 9 days.

(A1) The billing division is sending reminder emails to its oldest 25% of accounts. How many days old should an account at least be to receive a reminder?

(A2) Wildcat Electronics charges a late fee of $50 when an account is 30 days or older. What percentage of the accounts will receive this late fee?

(A3) Wildcat Electronics wants to give a discount to those accounts that pay their balance within 7 days. What percentage of the accounts will receive the discount?
Answer #1

1). we have

mean u=12 days

std deviation σ =9 days

a. oldest 25% means top 25%

so

P(X>x)=0.25 or

P(X<x)=0.75 or

P(Z<z)=0.75

z=0.68

so

0.68=(x-12)/9 or

x=18.12 days old,so it should be 18.12 days old to get a reminder.

b.

P(X>30)=P(Z>(30-12)/9)

=P(Z>2)=0.0228

so, 2.28% accounts will be charged with a fine of $50.

c.

within 7 days means

P(X<7)=P(Z<(7-12)/9)

=P(Z<-0.56)=0.2877

so, 28.77% accounts will be given a discount.

Please leave a like. Thank you.